An 29 g sample of an isotope of Strontium-90 was formed in 1960 explosion of an atomic bomb at

Johnson Island in the Pacific test Site. In what year will only 0.625 grams of the Strontium-90 remain?

Half-life of Strontium-90 is 28.8 years


1
Expert's answer
2021-09-18T23:57:56-0400

Half life 28.8years×2

0.625×29

= 28×2

56 years

= 56+1960= 2016